Critical angle of glass is  θ1 and the critical angle of water is θ2. Calculate the critical angle for water and glass surface. (μg = 3/2, μw = 4/3)

Grade:10

2 Answers

Prudhvi teja
13 years ago

Dear kaushik

The critical angle is the angle of incidence above which total internal reflection occurs. The angle of incidence is measured with respect to the normal at the refractive boundary. The critical angle θc is given by:

\theta_c = \arcsin \left( \frac{n_2}{n_1} \right),
critical angle = arcsin(4/3 / 3/2) = 62.73 degres



Please feel free to post as many doubts on our discussion forum as you can.we will get you the answer and detailed  solution very  quickly.


All the best.

Regards,
Askiitians Experts
Prudhvi Teja

 

 


Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian

arjun murali
11 years ago

can the answer be explained in any other proper way........cause i am not satisfied with that solution :(